OpenStudy (turingtest):

Prove that\[\int_{0}^{\pi}\frac{x\sin x}{1+\cos^2x}dx=\frac{\pi^2}{4}\]

OpenStudy (turingtest):

Thanks to Zarkon's tips I've got it up until\[-x\tan^{-1}(\cos x)+\int_{0}^{\pi}\tan^{-1}(\cos x)dx\]Which of course I don't know how to integrate. His tip from here was to use the bounds of integration and the fact that the arctangent is an odd function. Obviously the cosine is even, so arctan(cosx) is even as well I suppose, but he said there is a way to fix that. Still haven't figured out how to do that...

OpenStudy (zarkon):

you should really evaluate the first part from 0 to Pi

OpenStudy (turingtest):

\[\frac{\pi^2}{4}+\int_{0}^{\pi}\tan^{-1}(\cos x)dx\]Ok so it looks like we almost have the answer, but for proving that the integral is zero, as mentioned last night.

OpenStudy (zarkon):

use the hint i gave yesterday

OpenStudy (turingtest):

I woke up thinking about it I promise :)

OpenStudy (turingtest):

I mention above the tip you are referring to, right? Arctan odd, fix cosx somehow...

OpenStudy (zarkon):

yes...and that the integral of and odd function from -a to a is zero

OpenStudy (turingtest):

of course, I'm not that remedial...

OpenStudy (zarkon):

:)...sorry

OpenStudy (turingtest):

No, I'm sure it seems like it sometimes. but how to fix cosx...

OpenStudy (asnaseer):

does it involve using the half tangent formulas?

OpenStudy (mr.math):

I've got a much easier method to solve it.

OpenStudy (zarkon):

no..simpler :)

OpenStudy (asnaseer):

convert the integral from 0 to pi TO -pi/2 to pi/2?

OpenStudy (zarkon):

yes

OpenStudy (mr.math):

Lets substitute \(u=x+\frac{\pi}{2}\), then (call the integral I): \[I=\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\frac{(\frac{\pi}{2}-u)\cos(u)}{1+\sin^2u}=\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\frac{\frac{\pi}{2}\cos(u)}{1+\sin^2(u)}.\]

OpenStudy (mr.math):

That's because \(\frac{u\cos(u)}{1+\sin^2(u}\) is odd, so its integration from -pi/2 to pi/2 is 0. The rest is trivial.

OpenStudy (turingtest):

That's actually more like what I was trying to do, split up the integral, but I couldn't figure out how.

OpenStudy (zarkon):

I shifted the limits and use the fact that \[\cos(x+\pi/2)=-\sin(x)\] which is odd..the composition of odd functions is odd

OpenStudy (mr.math):

Typo, the substitution is \(u=\frac{\pi}{2}-x\). The rest is ok I think.

OpenStudy (turingtest):

Yes, I could see that was much simpler once asnaseer said it.

OpenStudy (mr.math):

Oh asnaseer already said that.

OpenStudy (asnaseer):

Zarkon is the one who left BIG FOOTPRINT clues all over the place - we just took a long time noticing them :-)

OpenStudy (turingtest):

It's so obvious once you know the answer XD

OpenStudy (zarkon):

a lot of problem are like that

OpenStudy (turingtest):

There's one more I'm gonna post from the calc section of the test. The rest is liner algebra. I don't think I stand a chance with that yet, need to brush up. Thanks again everybody!